# Pushing the Envelope of Low-Bit LLM via Dynamic Error Compensation

### v1

@@ 5.4 End-to-End Evaluation
Methodology. In this section, we present case studies demonstrating how QDEC advances model quality while minimizing latency increases across three desktop GPUs—RTX 4090, 4080 Super (4080S), and 4070 Super (4070S)—as well as two laptop GPUs—RTX 4070 Mobile (4070M), and 4050 Mobile (4050M). The specifications for these GPUs are listed in Table 2. We run the tuner for QDEC with four target slowdown rates (2.5%, 5%, 10%, 20%), and evaluate perplexity on WikiText as well as end-to-end inference latency with the resulting configurations. We integrate our kernel into a PyTorch-based inference pipeline optimized with the torch.compile feature [6] and measure the average time taken per token generation over 1024 tokens.
AWQ 3-bit SqueezeLLM 3-bit
GPU Target Tuner Results Slowdown Tuner Results Slowdown
4090 2.5% 24 / (4, 4, 8, 9) 1.3% 56 / (2, 0, 2, 2) 2.0%
5% 24 / (5, 7, 9, 10) 2.2% 56 / (1, 1, 2, 6) 4.0%
10% 24 / (10, 11, 11, 11) 4.9% 38 / (5, 4, 5, 7) 6.1%
20% 24 / (15, 15, 16, 15) 10.4% 28 / (10, 10, 10, 10) 10.8%
4080S 2.5% 6 / (9, 5, 9, 6) 1.1% 24 / (10, 11, 14, 16) 0.6%
5% 16 / (10, 10, 11, 9) 2.5% 24 / (14, 14, 14, 14) 2.2%
10% 24 / (18, 20, 23, 18) 5.5% 24 / (18, 19, 18, 17) 4.7%
20% 24 / (26, 27, 25, 24) 11.5% 38 / (23, 24, 22, 23) 10.8%
4070S 2.5% 28 / (25, 26, 26, 26) 2.0% 19 / (16, 15, 24, 21) 1.5%
5% 24 / (31, 31, 35, 29) 3.7% 24 / (21, 22, 24, 24) 3.0%
10% 24 / (35, 35, 36, 34) 7.1% 24 / (28, 30, 30, 27) 6.2%
20% 28 / (44, 44, 41, 42) 13.9% 24 / (35, 36, 34, 35) 12.4%
4070M 2.5% 16 / (38, 39, 39, 40) 1.7% 12 / (28, 28, 35, 29) 1.6%
5% 16 / (42, 43, 42, 42) 3.4% 16 / (33, 34, 36, 33) 3.1%
10% 16 / (46, 46, 44, 45) 6.6% 16 / (37, 38, 37, 37) 6.1%
20% 16 / (50, 50, 49, 50) 13.4% 16 / (42, 42, 41, 41) 12.7%
4050M 2.5% 8 / (55, 56, 58, 55) 1.7% 7 / (45, 45, 50, 44) 1.7%
5% 8 / (59, 59, 59, 58) 3.3% 9 / (48, 48, 48, 48) 3.3%
10% 10 / (62, 62, 62, 62) 6.6% 10 / (52, 53, 52, 52) 6.3%
20% 10 / (70, 70, 68, 68) 13.5% 10 / (59, 60, 57, 58) 12.6%
Table 3: Tuner results ($n^{max}_{tb}/(k^{qkv}_{chunk},k^{o}_{chunk},k^{gu}_{chunk},k^{d}_{chunk})$) for four target slowdown rates (2.5%, 5%, 10%, 20%) and corresponding actual slowdown rates for 3-bit Llama-3.
@@
Figure 18: Perplexity against time per token on various NVIDIA GPUs. The x markers indicate baseline values, where QDEC is not applied ($k_{chunk}=0$). Subsequent markers show DecDEC results on target slowdown rates 2.5%, 5%, 10%, and 20%.
@@
Results. Table 3 lists the configurations obtained from the tuner along with the actual end-to-end slowdowns compared to the baseline. Due to space limits, we report only the results for 3-bit Llama-3. Similar trends are observed for the 4-bit configurations, as well as for Phi-3. In all cases, the actual slowdown is below the target rate. This is expected as the tuner configures parameters conservatively. The tuner targets only the kernel times of linear operations, while other operations outside the linear layers (e.g., attention, normalization) also contribute to overall inference time. Consistent with the observations in Section 5.3, in general, selected $k$ values are higher for GPUs with greater PCIe-to-GPU memory bandwidth ratio (4050M $>$ 4070M $\simeq$ 4070S $>$ 4080S $>$ 4090).
@@
Figure 18 shows the trends of end-to-end inference latency versus perplexity. All Phi-3 cases, as well as AWQ 3.5-bit, 4-bit, and SqueezeLLM 4-bit cases of Llama-3 face out-of-memory issues on the 4050M, and are thus excluded. Similarly, the AWQ 4-bit case of Phi-3 for 4070M is also excluded. In each line in Figure 18, the x marker represents the baseline, while subsequent markers show the results for QDEC on the four target slowdown rates, in increasing order.
@@
For all cases, QDEC demonstrates promising Pareto-optimal trade-offs between model quality and inference latency. The results for target slowdown rates of 2.5% and 5% are particularly impressive, while cases with 10% and 20% slowdown rates tend to exhibit diminishing returns in perplexity reduction. On platforms with high PCIe-to-GPU memory bandwidth ratios, such as the 4070S, 4070M, and 4050M, incorporating QDEC with a 2.5% target slowdown rate on 3-bit models occasionally results in lower perplexity than their 3.5-bit counterparts. In these scenarios, QDEC provides Pareto-dominant solutions by excelling in model quality, latency, and GPU memory consumption. Examples include AWQ Llama-3 on 4070M and 4050M, as well as AWQ Phi-3 on 4070S and 4070M. A particularly noteworthy case is the AWQ 3-bit Llama-3 on 4050M, where QDEC reduces perplexity from 10.15 to 9.12 with only a 1.7% latency slowdown (highlighted by a red circle). This perplexity is lower than the 3.5-bit baseline, which has been unattainable on this platform without QDEC due to memory capacity constraints. This case highlights QDEC’s effectiveness in pushing the boundaries of quantized LLMs within memory capacity constraints.
@@ 6 Related Work
GPU Implementations for Weight-Only Quantization. Numerous studies have proposed efficient GPU implementations tailored for weight-only quantization of LLMs. LUTGEMM [44] replaces the expensive dequantization with simple LUT operations. Marlin [20] supports FP16-INT4 GEMM across various batch sizes, and Quant-LLM [62] introduces a GPU kernel that efficiently handles non-power-of-2 bitwidths (e.g., 6-bit). FLUTE [23] provides a specialized kernel for non-uniform quantization. Any-Precision LLM [45] suggests a memory-efficient kernel for adaptive bitwidth selection. These implementations can seamlessly integrate with QDEC, benefiting from its dynamic error compensation.
LLM Inference with External Memory. Several works address GPU memory limitations in LLM inference by leveraging external memories like CPU memory or disk. DeepSpeed-Inference [5] and FlexGen [49] focus on throughput-oriented, out-of-core inference. Pre-gated MoE [28] retrieves the parameters of only the activated experts in MoE models from CPU memory. LLM-in-a-Flash [4] introduces a flash memory-based inference system exploiting activation sparsity. InfiniGen [34] offloads the key-value cache to CPU memory. PowerInfer [50] proposes a GPU-CPU hybrid inference engine that leverages activation sparsity in ReLU-based models. Though these approaches share the same goal as QDEC in aiming to extend GPU memory, they address distinct scenarios with different challenges and opportunities.
LLM Compression Methods. In addition to quantization, pruning can improve the efficiency of LLM inference [18, 52, 37]. Other techniques, including knowledge distillation [25, 22] and low-rank decomposition [66, 47], offer additional ways to compress LLMs. These methods are orthogonal to quantization and, consequently, to QDEC.
@@ 7 Conclusion
We propose QDEC, an inference scheme for low-bit LLMs that improves model quality by correcting quantization errors through selective retrieval of residuals stored in CPU memory. By focusing on dynamically identified salient channels at each decoding step, QDEC maximizes error compensation within a limited transfer volume. QDEC significantly improves quality of quantized LLMs with minimal overheads.
Appendix A Identifying Viable Values for $n_{tb}$ and $k_{chunk}$
A.1 Identifying Viable Values for $n_{tb}$
Viable $n_{tb}$ values for each linear layer are restricted to a finite set $S$, determined by the input and output channel numbers ($c_{in}$ and $c_{out}$). The QDEC kernel comprises two phases—channel selection and residual fetch with GEMV—each of which places constraints on how thread blocks can be effectively utilized.
An $n_{tb}$ value is considered non-viable if it results in redundant thread blocks during both the channel selection and residual fetch with GEMV phases, as such redundancy persists throughout the entire QDEC kernel. Therefore, for the set of usable thread block numbers in channel selection ($A$) and the set of usable thread block numbers in residual fetch with GEMV ($B$), the set of viable $n_{tb}$ values ($N$) is given by:
$N=A\cup B.$
Constraints in Channel Selection. QDEC employs a chunked approximate Top-K method for efficiently identifying salient channels. During the approximate Top-K process, synchronization is required to track the number of selected channels within each chunk. To achieve this, each chunk is assigned to a single thread block, avoiding the need for inter-thread block communication or global memory synchronization, which would otherwise be necessary.
As a result, the number of thread blocks that can be used is constrained by the number of chunks. However, it is also possible to use fewer thread blocks, as a single thread block can manage multiple chunks. Therefore, the set of usable thread block numbers in channel selection, $A$, is defined as:
$A=\{n\mid n\in\mathbb{Z},1\leq n\leq\frac{c_{in}}{\text{(chunk size)}}\}.$
Constraints in Residual Fetch & GEMV. The residual fetch & GEMV phase employs CUDA zero-copy operations to retrieve quantized residuals over PCIe. The 4-bit quantized residuals are packed into 32-bit integers, with each integer containing 8 quantized values. Each thread reads one 32-bit integer via zero-copy, resulting in 8 quantized values per thread. To minimize memory transfer overheads, zero-copy requests are coalesced. Given that a single warp consists of 32 threads, a coalesced zero-copy request transfers 256 quantized values in a 128-byte packet.
Since QDEC cannot determine in advance which input channels will be selected for fetching, the residuals are stored with output channels arranged contiguously in memory. For each selected input channel, QDEC fetches the residuals corresponding to all output channels. These residuals, equal to the number of output channels, are retrieved in contiguous segments of 256 elements to ensure coalesced access.
The total number of segments to be fetched, $s$, is given by $s=\frac{c_{out}}{256}$. These segments must be distributed among $n$ thread blocks, so the workload per thread block is defined as $\lceil\frac{s}{n}\rceil$, where the ceiling operation ensures that no segments are omitted. However, for some values of $n$, this distribution can result in redundant thread blocks that do not process any segments.
For instance, consider $n=7$ and $c_{out}=6144$. The total number of segments is $s=\frac{c_{out}}{256}=24$. With $n=7$, each thread block is assigned $\lceil\frac{24}{7}\rceil=4$ segments. However, since each thread block handles 4 segments, only 6 thread blocks are needed to process all 24 segments. This raises the question of whether redistributing the workload could justify the use of the $n=7$ configuration. For example, one could attempt to have 4 thread blocks process 3 segments each, while the remaining 3 thread blocks process 4 segments each (totaling 24 segments). However, this does not improve efficiency, as the execution time remains bottlenecked by the thread blocks processing 4 segments. Thus, $n=7$ does not provide any advantage over $n=6$ and is a redundant configuration.
To avoid such inefficiency and cut down on the search space, we consider a thread block configuration viable only if no thread blocks become redundant when each thread block fetches $\lceil\frac{s}{n}\rceil$ segments. Specifically, $n$ is considered viable if:
$\left\lceil s\div\left\lceil\frac{s}{n}\right\rceil\right\rceil=n.$
This condition ensures that the workload distribution, defined by $\left\lceil\frac{s}{n}\right\rceil$ segments per thread block, cannot be achieved with fewer thread blocks, thereby eliminating redundant configurations.
The set of usable thread block numbers in residual fetch & GEMV, $B$, is thus defined as:
$B=\{n\mid n\in\mathbb{Z},1\leq n\leq s,\left\lceil s\div\left\lceil\frac{s}{n}\right\rceil\right\rceil=n,s=\frac{c_{out}}{256}\}.$
Memory Constraints and Practical Adjustments. In practice, the number of segments processed by each thread block may be constrained by hardware resources, such as the available shared memory. For certain smaller $n$ values and larger $c_{out}$, the number of segments to fetch per thread block can exceed the shared memory capacity (see Section A.2 for exact details). In such cases, QDEC reduces the workload by splitting the work over the output channels into multiple passes (e.g., processing half the output channels in each pass and repeating twice). This adjustment allows QDEC to operate within hardware constraints but has the side effect of effectively reducing $c_{out}$, and consequently $s$, for those specific configurations. These changes can alter the set of viable $n$ values, as the workload distribution and associated constraints are modified.
Example: Viable $n_{tb}$ Calculation. Consider a scenario where $c_{in}=4096$, $c_{out}=6144$, and the chunk size is set to 1024.
$A$ is the set of usable thread block numbers for channel selection, as previously defined:
$A$ =$=\{n\mid n\in\mathbb{Z},1\leq n\leq\frac{c_{in}}{\text{chunk size}}\}$
=$=\{n\mid n\in\mathbb{Z},1\leq n\leq\frac{4096}{1024}\}$
=$=\{n\mid n\in\mathbb{Z},1\leq n\leq 4\}.$
$B$ is the set of usable thread block numbers for residual fetch with GEMV, as previously defined:
$\begin{split}B&=\{n\mid n\in\mathbb{Z},1\leq n\leq s,\left\lceil s\div\left\lceil\frac{s}{n}\right\rceil\right\rceil=n,s=\frac{c_{out}}{256}\}\\ &=\{n\mid n\in\mathbb{Z},1\leq n\leq 24,\left\lceil 24\div\left\lceil\frac{24}{n}\right\rceil\right\rceil=n\}\\ &=\{1,2,3,4,5,6,8,12,24\}.\end{split}$
The final set of viable $n_{tb}$ values is:
$N=A\cup B=\{1,2,3,4,5,6,8,12,24\}.$
Summary and Practical Implications. The constraints on $n_{tb}$ values for the QDEC kernel arise from the requirements of both the channel selection and residual fetch with GEMV phases. In channel selection, $n_{tb}$ is constrained by the chunk structure and synchronization needs, while in residual fetch with GEMV, it is determined by the need for coalesced memory access and the elimination of redundant thread blocks. By integrating these constraints, the set of viable $n_{tb}$ values, $N=A\cup B$, ensures thread block configurations that maximize computational efficiency and memory access performance.
In practice, users can set $n^{max}_{tb}$, an upper limit on the number of thread blocks to use in QDEC. Based on this limit, QDEC automatically selects the largest viable $n_{tb}$ value from $N$, balancing performance with the constraints of the hardware and kernel design.
A.2 Identifying Viable Values for $k_{chunk}$
The maximum value of $k_{chunk}$ is constrained by the shared memory available per block, a resource that varies across GPU architectures. In the GPUs tested in this study, the shared memory per block is limited to 48 KB. Both statically and dynamically allocated shared memory contribute to the total usage, and this combined usage determines the upper bound for $k_{chunk}$.
Shared Memory Usage in QDEC. The shared memory in QDEC is divided into static and dynamic components. Static shared memory usage is proportional to the number of buckets used in the approximate Top-K process and is fixed at 128 bytes for 32 buckets.
Dynamic shared memory is allocated for the channel selection and residual fetch & GEMV phases. Since memory is recycled between these two phases, the total shared memory required is determined by the phase with the higher memory demand.
For the channel selection phase, the dynamic shared memory usage is:
[…]
For the residual fetch & GEMV phase, the dynamic shared memory usage is:
$6\times k+16384+512\times\left\lceil\frac{s}{n_{tb}}\right\rceil\text{ bytes},$
where $s=\frac{c_{out}}{256}$ as defined in Section A.1. If the memory requirement for this phase exceeds the available shared memory, a fallback mode reduces memory usage to:
$16384+512\times\left\lceil\frac{s}{n_{tb}}\right\rceil\text{ bytes}.$
This fallback mode trades reduced memory usage for a slight increase in kernel runtime.
Once the dominant dynamic memory usage is identified, the static memory is added to calculate the total shared memory required.
Example: Calculating the Upper Bound of $k_{chunk}$. To illustrate how shared memory limits the maximum $k_{chunk}$, consider a scenario where $c_{in}=4096$, $c_{out}=6144$, and the chunk size is set to 1024.
The total number of chunks is:
$\frac{c_{in}}{\text{chunk size}}=\frac{4096}{1024}=4.$
Since $k=4\times k_{chunk}$, the memory requirement for the channel selection phase is:
$128\times k_{chunk}+2\times 1024\text{ bytes}.$
The memory requirement for the residual fetch & GEMV phase is:
$6\times k+16384+512\times\left\lceil\frac{24}{n_{tb}}\right\rceil\text{ bytes}.$
In this example, the channel selection phase requires more memory as $k_{chunk}$ increases, due to the higher coefficient. Therefore, the total shared memory required, including static shared memory, is:
$128\times k_{chunk}+2048+128\text{ bytes}.$
On a GPU with 48 KB (49,152 bytes) of shared memory per block, the maximum $k_{chunk}$ can be calculated as:
$k_{chunk}=\frac{49152-128-2048}{128}=367.$
Thus, for the given parameters, the maximum $k_{chunk}$ is 367. Beyond this value, the shared memory requirements exceed the hardware limit.

### v2

@@ 5.3 End-to-End Evaluation
Methodology. In this section, we present case studies demonstrating how DecDEC advances model quality while minimizing latency increases across three desktop GPUs—RTX 4090, 4080 Super (4080S), and 4070 Super (4070S)—as well as two laptop GPUs—RTX 4070 Mobile (4070M), and 4050 Mobile (4050M). The specifications for these GPUs are listed in Table 1. We run the tuner for DecDEC with four target slowdown rates (2.5%, 5%, 10%, 20%), and evaluate perplexity on WikiText as well as end-to-end inference latency with the resulting configurations. For the base GEMV kernel, we use LUTGEMM [44] for AWQ and Any-Precision LLM [45] for SqueezeLLM—both state-of-the-art kernels for uniform and non-uniform quantization, respectively. We integrate DecDEC into a PyTorch-based inference pipeline optimized with the torch.compile feature [5] and measure the average time taken per token generation over 1024 tokens.
For the 3.5-bit models, we do not run the tuner separately. Instead, we construct the configuration by combining the tuning results with the 3-bit and 4-bit models. Specifically, we use the configuration from tuning with the 3-bit model for layers quantized to 3 bits, and the configuration from tuning with the 4-bit model for layers quantized to 4 bits—both assuming the same target slowdown rate.
@@
Figure 17: Perplexity against time per token on various NVIDIA GPUs. The x markers indicate baseline values, where DecDEC is not applied ($k_{chunk}=0$). Subsequent markers show DecDEC results on target slowdown rates 2.5%, 5%, 10%, and 20%.
Llama-3-8B-Instruct Phi-3-medium-4k-instruct
AWQ 3-bit SqueezeLLM 3-bit AWQ 3-bit SqueezeLLM 3-bit
GPU Target Tuner Results Slowdown Tuner Results Slowdown Tuner Results Slowdown Tuner Results Slowdown
4090 2.5% 24 / (4, 4, 8, 9) 1.3% 56 / (2, 0, 2, 2) 2.0% 20 / (8, 6, 13, 13) 1.5% 47 / (5, 2, 7, 10) 1.8%
5% 24 / (5, 7, 9, 10) 2.2% 56 / (1, 1, 2, 6) 4.0% 47 / (10, 11, 13, 13) 3.8% 35 / (6, 6, 9, 10) 3.5%
10% 24 / (10, 11, 11, 11) 4.9% 38 / (5, 4, 5, 7) 6.1% 35 / (15, 15, 15, 14) 7.8% 47 / (10, 11, 13, 11) 7.3%
20% 24 / (15, 15, 16, 15) 10.4% 28 / (10, 10, 10, 10) 10.8% 20 / (18, 19, 18, 18) 15.0% 35 / (16, 16, 15, 14) 15.3%
4080S 2.5% 6 / (9, 5, 9, 6) 1.1% 24 / (10, 11, 14, 16) 0.6% 30 / (18, 17, 18, 20) 1.7% 30 / (14, 14, 15, 16) 2.0%
5% 16 / (10, 10, 11, 9) 2.5% 24 / (14, 14, 14, 14) 2.2% 30 / (21, 21, 23, 21) 3.9% 30 / (17, 17, 20, 19) 3.7%
10% 24 / (18, 20, 23, 18) 5.5% 24 / (18, 19, 18, 17) 4.7% 35 / (24, 25, 24, 24) 8.3% 35 / (22, 22, 23, 21) 7.6%
20% 24 / (26, 27, 25, 24) 11.5% 38 / (23, 24, 22, 23) 10.8% 35 / (30, 32, 28, 29) 16.2% 35 / (27, 27, 25, 25) 15.1%
4070S 2.5% 28 / (25, 26, 26, 26) 2.0% 19 / (16, 15, 24, 21) 1.5% 20 / (31, 31, 37, 32) 2.1% 20 / (25, 26, 30, 28) 2.1%
5% 24 / (31, 31, 35, 29) 3.7% 24 / (21, 22, 24, 24) 3.0% 20 / (35, 35, 37, 34) 4.2% 20 / (30, 30, 30, 29) 4.2%
10% 24 / (35, 35, 36, 34) 7.1% 24 / (28, 30, 30, 27) 6.2% 24 / (39, 40, 41, 38) 8.5% 24 / (34, 35, 35, 33) 8.6%
20% 28 / (44, 44, 41, 42) 13.9% 24 / (35, 36, 34, 35) 12.4% 20 / (45, 46, 44, 45) 17.0% 20 / (40, 41, 38, 39) 16.9%
4070M 2.5% 16 / (38, 39, 39, 40) 1.7% 12 / (28, 28, 35, 29) 1.6% 18 / (41, 41, 41, 42) 2.1% 17 / (33, 35, 35, 34) 2.1%
5% 16 / (42, 43, 42, 42) 3.4% 16 / (33, 34, 36, 33) 3.1% 18 / (44, 46, 43, 44) 4.4% 18 / (37, 38, 37, 37) 4.3%
10% 16 / (46, 46, 44, 45) 6.6% 16 / (37, 38, 37, 37) 6.1% 18 / (46, 48, 46, 46) 8.3% 18 / (40, 41, 39, 39) 8.5%
20% 16 / (50, 50, 49, 50) 13.4% 16 / (42, 42, 41, 41) 12.7% 18 / (53, 53, 50, 51) 17.2% 18 / (44, 44, 43, 43) 16.7%
4050M 2.5% 8 / (55, 56, 58, 55) 1.7% 7 / (45, 45, 50, 44) 1.7% OOM - OOM -
5% 8 / (59, 59, 59, 58) 3.3% 9 / (48, 48, 48, 48) 3.3% OOM - OOM -
10% 10 / (62, 62, 62, 62) 6.6% 10 / (52, 53, 52, 52) 6.3% OOM - OOM -
20% 10 / (70, 70, 68, 68) 13.5% 10 / (59, 60, 57, 58) 12.6% OOM - OOM -
Table 3: Tuner results ($n^{max}_{tb}/(k^{qkv}_{chunk},k^{o}_{chunk},k^{gu}_{chunk},k^{d}_{chunk})$) for four target slowdown rates (2.5%, 5%, 10%, 20%) and corresponding actual slowdown rates for 3-bit Llama-3 and Phi-3. Phi-3 is out of memory (OOM) on the 4050M GPU.
@@
Results. Table 3 lists the configurations obtained from the tuner along with the actual end-to-end slowdowns compared to the baseline. We report only the results for 3-bit as similar trends are observed for the 3.5-bit and 4-bit configurations. In all cases, the actual slowdown is below the target rate. This is expected as the tuner configures parameters conservatively. The tuner targets only the kernel times of linear operations, while other operations outside the linear layers (e.g., attention, normalization) also contribute to overall inference time. Consistent with the observations in Section 5.1, in general, selected $k$ values are higher for GPUs with greater PCIe-to-GPU memory bandwidth ratio (4050M $>$ 4070M $\simeq$ 4070S $>$ 4080S $>$ 4090).
@@
Figure 17 shows the trends of end-to-end inference latency versus perplexity. All Phi-3 cases, as well as AWQ 3.5-bit, 4-bit, and SqueezeLLM 4-bit cases of Llama-3 face out-of-memory issues on the 4050M, and are thus excluded. Similarly, the AWQ 4-bit case of Phi-3 for 4070M is also excluded. In each line in Figure 17, the x marker represents the baseline, while subsequent markers show the results for DecDEC on the four target slowdown rates, in increasing order.
@@
For all cases, DecDEC demonstrates promising Pareto-optimal trade-offs between model quality and inference latency. The results for target slowdown rates of 2.5% and 5% are particularly impressive, while 10%–20% show diminishing returns. On platforms with high PCIe-to-GPU memory bandwidth ratios, such as the 4070S, 4070M, and 4050M, DecDEC at 2.5% slowdown on 3-bit models sometimes outperforms 3.5-bit baselines. In these cases, DecDEC yields Pareto-dominant solutions by excelling in model quality, latency, and memory. Examples include AWQ Llama-3 (4070M, 4050M) and AWQ Phi-3 (4070S, 4070M). A particularly noteworthy case is the AWQ 3-bit Llama-3 on 4050M, where DecDEC reduces perplexity from 10.15 to 9.12 with only a 1.7% latency slowdown (highlighted by a red circle). This outperforms the 3.5-bit baseline, which is infeasible on 4050M without DecDEC due to memory limits. This case highlights DecDEC’s effectiveness in pushing the boundaries of quantized LLMs within memory capacity constraints.
Figure 18: Perplexity vs. time per token: (a) across different GPU generations and (b) on server-grade GPUs.
5.4 Evaluation across GPU Generations
We evaluate the robustness of DecDEC across GPU generations, as both GPU memory bandwidth and PCIe bandwidth, two key factors influencing its effectiveness, may increase with newer architectures. To this end, we use three consumer-grade GPUs from the same product class across different generations—RTX 3080, 4080S, and 5080—whose specifications are listed in Table 4. The $R_{bw}$ values (lower is better) remain essentially unchanged from the 3080 to the 4080S and even decrease from the 4080S to the 5080. Figure 18(a) plots the perplexity of AWQ-quantized Phi-3 with DecDEC augmentation against end-to-end latency on these GPUs, using the methodology described in Section 5.3. The improvements delivered by DecDEC are comparable across all three cards, confirming its robustness across generations.
5.5 Applicability to Server-Grade GPUs
Although DecDEC primarily targets single-batch inference on client (edge) devices, this section evaluates its effectiveness on server-grade GPUs. We measure perplexity versus end-to-end latency using an AWQ-quantized Llama-3-70B-Instruct model augmented with DecDEC on the H100 SMX5 and GH200, following the methodology described in Section 5.3. Both GPUs provide 3.36 TB/s of memory bandwidth; however, the GH200’s 450 GB/s NVLink-C2C far exceeds the H100’s 64 GB/s PCIe, yielding a much lower $R_{bw}$.
Figure 18(b) shows the results. DecDEC improves perplexity on both GPUs with minimal latency overhead. However, the GH200’s advantage is smaller than the $R_{bw}$ gap suggests. DecDEC assumes a DRAM-bound GEMV, where reallocating SMs for error compensation do not impact GEMV latency. On server GPUs, however, quantized GEMV (e.g., LUTGEMM [44]) is L1-bound, not DRAM-bound. Since L1 throughput scales with active SMs, reallocating SMs increases GEMV latency, limiting DecDEC’s benefit despite low $R_{bw}$. Enhancing quantized GEMV kernels for server-grade GPUs by mitigating L1 bottlenecks could unlock further gains.
Memory Bandwidth PCIe Bandwidth $\bm{R_{bw}}$
RTX 5080 960 GB/s 64 GB/s 15
RTX 4080S 736 GB/s 32 GB/s 23
RTX 3080 760 GB/s 32 GB/s 24
Table 4: 80-class GPU specifications across generations.
@@ 6 Related Work
GPU Implementations for Weight-Only Quantization. Numerous studies have proposed efficient GPU implementations tailored for weight-only quantization of LLMs. LUTGEMM [44] replaces the expensive dequantization with simple LUT operations. Marlin [20] supports FP16-INT4 GEMM across various batch sizes, and Quant-LLM [63] introduces a GPU kernel that efficiently handles non-power-of-2 bitwidths (e.g., 6-bit). FLUTE [23] provides a specialized kernel for non-uniform quantization. Any-Precision LLM [45] suggests a memory-efficient kernel for adaptive bitwidth selection. These implementations can seamlessly integrate with DecDEC, benefiting from its dynamic error compensation.
LLM Inference with External Memory. Several works address GPU memory limitations in LLM inference by leveraging external memories like CPU memory or disk. DeepSpeed-Inference [4] and FlexGen [50] focus on throughput-oriented, out-of-core LLM inference. Pre-gated MoE [28] retrieves the parameters of only the activated experts in MoE models from CPU memory. LLM-in-a-Flash [3] introduces a flash memory-based inference system exploiting activation sparsity. InfiniGen [34] offloads the key-value cache to CPU memory. PowerInfer [51] proposes a GPU-CPU hybrid inference engine that leverages activation sparsity in ReLU-based models. Though these approaches share the same goal as DecDEC in aiming to extend GPU memory, they address distinct scenarios with different challenges and opportunities.
LLM Compression Methods. In addition to quantization, pruning can improve the efficiency of LLM inference [18, 53, 37]. Other techniques, including knowledge distillation [25, 22] and low-rank decomposition [67, 48], offer additional ways to compress LLMs. These methods are orthogonal to quantization and, consequently, to DecDEC.
@@ 7 Conclusion
We propose DecDEC, an inference scheme for low-bit LLMs that improves model quality by correcting quantization errors through selective retrieval of residuals stored in CPU memory. By focusing on dynamically identified salient channels at each decoding step, DecDEC maximizes error compensation within a limited transfer volume. DecDEC significantly improves quality of quantized LLMs with minimal overheads.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) (RS-2024-00340008, RS-2024-00405857). Jae W. Lee is the corresponding author.